Suppose consumer can purchase food (good 1) or clothing (good 2) at prices p1 = $5 and p2 = $5 per unit with her income of m = $100. The government gives her an endowment of 4 food stamps where each stamp entitles her to one unit of food. We will assume that food, clothing and food stamps are divisible. Draw his budget line.
Answer #1

Clothing 20 - Budget line - 24 Food

Y-intercept = m/P2 = 100/5 =20

X -intercept = 4+m/p1 = 4+20 = 24